Suppose 100 million people are in the civilian labor force and 90 7. million people are employed. How many people are unemployed? What is the unemployment rate?
1
Expert's answer
2020-02-13T08:37:38-0500

unemployed, U= 100-90.7= 9.3million

unemployment rate, "UR= (U\/Lf )*100" =( 9.3/100) *100 = 9.3%